1 unit of Currency A costs 1.47 units of Currency B, 1 unit of Currency B costs 0.65 units of Currency C, and 1 unit of Currency C costs 0.98 units of Currency A. Assume that you are starting out with 1,000,000 units of Currency A and that there are no transaction costs. How much money will you make in one set of triangular arbitrage transactions? Round to the nearest unit of Currency A.
Answer #1

Convert Currency A into Currency C and get 1,000,000/0.98 = 1,020,408.16

Convert into Currency B and get 1,020,408.16/0.65 = 1,569,858.71 units

Convert into Currency A and get 1,569,858.71/1.47 = 1,067,931.10

Hence, money earned from triangular arbitrage = 1,067,931 - 1,000,000 = 67,931 units